Prev Next

SW&IT Meeting Minutes 2018-11-26

Attendees: Andrew, David, Matthieu, Petr, Stephen, Yang

Outlook

Labor Reporting

Report auto-generated every, Do not edit this table, use entries in smartsheets. 20181105
LR20181105 - Bec, Matthieu 81.0
LR20181105 - Colson, Andrew 40.0
LR20181105 - Edmeades, David 32.0
LR20181105 - Sola, Igor 81.0
LR20181105 - Kubanek, Petr 65.5
LR20181105 - Hooper, Stephen 80.0
LR20181105 - Zhang, Yang 40.0
LR20181105 - Zhang, Xianyu 72.0

20181119
LR20181119 - Bec, Matthieu 37.0
LR20181119 - Colson, Andrew 0.0
LR20181119 - Edmeades, David 24.0
LR20181119 - Sola, Igor 25.0
LR20181119 - Kubanek, Petr 18.0
LR20181119 - Hooper, Stephen 64.0
LR20181119 - Zhang, Yang 0.0
LR20181119 - Zhang, Xianyu 0.0

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s)Sorted ascending Primary Secondary Week
ALH Igor(2)
Matthieu (2)
1
AzCam
AC (2)
Petr (2)
2
DIMM
AC (5)
Petr (2)
2
FACSUM,WEATHERSTATION
AC (2)
Matthieu (2)
4
GCS
Petr (5)
Matthieu (2)
1
INDI Igor (2)
Matthieu (2)
3
LBC
Yang (5)
Igor(2) 2
MODS
Yang (3)
Igor (1/2) 2
OAC
AC (5)
Petr (4)
1
OVMS
Petr (5)
Matthieu (2)
3
particle monitors AC Petr 4
Status Server Stephen (2)
Matthieu (2)
2
Telem-VIZ
Yang (1)
Stephen (1) 3
Telemetry
Petr (5)
Matthieu (2)
2
Telemetry cleanup Stephen (1) Petr 3
Weatherstation
AC (2)
Petr (2)
2

News from management

  • PEPSI/pol high profile run in January: AGW3 & 4 must be ready
  • on-site interview Wed

Previous Action Items

  • Skill set list (DE)

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • network dropout
  • revert simbad harvard-strasbourg /et/hosts hack from Friday
  • Dust monitor status
  • AO cabinet switch upgrade - coordinate wth Mountain and AO
  • wifi summit upgrade (note: end-of-year deadline)
  • Tucson CR power overhaul plan (note: end-of-year deadline)
  • n507/obs upgrade (note: end-of-year deadline)
  • call out phone list (AC cell)
  • Wiki formatting

This week in github

Team reports

PK

  • fixed position angle issue (IT7463), including patching a patch
  • almost produced cross compiling gcc running on Intel for PowerPC (OVMS)
next:

AC

  • ARGOS engineering nights
  • Bandwith tests over VNC and SSH
  • SAS-MCSPU communication
  • Add check for NaN in pyramid wfs slopes
  • Add a warning for the operator when TCS Data Dictionary is not available

YZ

  • Back from vacation.
  • Continue to work on Polymer/JS9 problems.
  • Continue to work on LBC/MODS documentation.

MB

  • OC, WPC, management
  • OVMS+TT follow up (AO cabinet space, IPC, aosupervisor changes, f/w etc.)
  • TMS closure
  • OCS planning
  • Js9 play, looks great
  • review DE power plan
  • delta-tau training
  • svn2git sciops
  • svn: vs git:MODS, contacted Rick and Jerry
  • alternatives to code-hilight in wiki
  • f29 play + download
next:
  • on-site interview
  • automate weekly report heading section
  • revamp software wiki main page
  • science archive updates from IA2
  • delta-tau training
  • f29 migration

ISS

Last Week:
  • LUCI
    • Real Time Display (RTD) not displaying AFC windows.
      • I think it's related to Aladin and Journalizer issues.
    • Added new AFC subwindow coordinates for new AFC longslit masks. Improve ingestion script and also figure out use of group in AFC spots.
    • Look into LUCI use of rotator type, currently only uses 'POSITION", "PARALLACTIC", "NATIVE"
  • OT
    • Target catalog search not failing over to second catalog choice.
This week:
  • LUCI
    • Real Time Display (RTD) not displaying AFC windows.
    • See if I can reproduce CompensationMirror daemon hang up with new Dave suggestion. (I.e. use SSR mode before selecting AFC option)
  • OT
    • Target catalog search not failing over to second catalog choice.

DE

  • Wrangled with purchasing over drives; PO finally issued and sent to vendor
  • Moving directories off of volume2. No reported problems re: business directory
  • AzCam babysitting
    • Handling the belt-and-suspenders DB15 connector anchoring on capture cards
  • Repository permission problem. No response from Cristina, but Stephen and I think we have solved it by changing the top level directory permissions. We will see this evening when the new directory is created
    • I manually fixed permissions on previous directories and files; everyone should have access
  • Worked on power plan; focusing on it Monday.
  • Drives unlikely to arrive this week, giving us time for final prep of the volumes for deletion when they do.

SH:

  • dockerized the display a display runner -- problem, no chromium for 32-bit.
  • DNS mostly done.
  • cleaned up a few cron jobs.
  • added some stuff to http://info
  • switched FosWiki over to Apache.
next:
  • continue working on Wifi summary
  • continue adding information to http://info
  • direct rsync between jobs.
  • upgrade to Fedora 29 on hosts -- yumsrv, cisrv01 and mil.tucson
  • changes to AO cabinet switches.
Topic revision: r16 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback